I have a question. I have to create an equation to show how long it would take a submarine to reach a deoth of 428metres if it goes down 50 m every 5 minutes. Thanks!
Answer by checkley77(12844)

428/50=X/5
50X=428*5
50X=2,140
X=2,140/50
X=42.8 MINUTES.